5. Consider the data points (x, y) (10,0.8), (20, 1.6), (30, 2.3), (40, 2.8) (with equal error bars ơi-0-0.06), which have been fitted by the straight line y-0.2 + 0.067x using a least squares method. (a) (4 pts.) Find the value of x2 for the data. (b) (4 pts) Is it possible to assert, with 95% confidence, that a straight line is a good fit for the data? Explain why or why not Hint: Once you have set up the required matrices, you may use a calculator] for the matrix operations, just mention where you used a calculator. 1If you dont have a calculator that handles matrices, a website like www.wolframalpha.com may help. either the instructor nor the T.A. will provide assistance in using calculators, software, or websites.
Answer #1

here we use chi-square test and chi-square=Sum((O-E)2/E)=0.0139 with k-1=4-1=3 df

with null hypothesis H0: observed value =expected value ( line is good fit)

alternate hypothesis Ha: observed value not =expected value ( line is not good fit)

(a) chi-square=0.0139

(b) the critical chi-square(0.05,3)= 7.8147 is the more than calculated chi-square=0.0139, so we fail to reject H0 and conclude that straight line is good fit

following calcualtion has been done for answering this question

x y(observed) y(expected) O-E (O-E)2/E
10 0.8 0.87 -0.07 0.0056
20 1.6 1.54 0.06 0.0023
30 2.3 2.21 0.09 0.0037
40 2.8 2.88 -0.08 0.0022
sum 7.5 7.5 -1.1E-15 0.0139

O=observed

E=Expected